given a+b=c+d and a3+b3=c3+d3


prove that a2011+b2011=c2011+d2011


plzz xplain

Grade:10

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
given a+b=c+d and a3+b3=c3+d3
So (a+b)(a2-ab+b2)=(c+d)(c2-cd+d2)
As a+b=c+d ;(a2-ab+b2)=(c2-cd+d2)
So (a+b)2-3ab=(c+d)2-3cd
So ab=cd
also (a+b)2-4ab=(c+d)2-4cd
So (a-b)2=(c-d)2
SO a+b=c+d and a-b=c-d
So a=c,b=d
Soa2011+b2011=c2011+d2011